In a large plant population that reproduces by pollinator-assisted random mating (outcrossing), the frequencies of the genotypes GG, Gg and gg are 0.04, 0.32 and 0.64, respectively. Honey bees are under threat of decline. A loss of the bee pollinator forces the population to reproduce exclusively by self-fertilization. What will happen to the frequencies of the homozygous and heterozygous genotypes in this population after one generation of self-fertilization? If the same conditions continue to exist for many generations, what will eventually happen to the genotype frequencies at this locus?
Answer #1

Allele frequency of G = 4+16/100 = 0.20

Allele frequency of g = 16+64/100 = 0.80

Self fertilization of GG = GG*GG

G(0.2) G(0.2)
G(0.2) GG (0.04) GG(0.04)
G(0.2) GG (0.04) GG(0.04)

Frequency of GG = 0.04+0.04+0.04+0.04

= 0.16

Increase in homozygous freqency

Self fertilization of gg= gg*gg

g(0.8) g(0.8)
g(0.8) gg(0.64) gg(0.64)
g(0.8) gg(0.64) gg(0.64)

frequency of gg = 0.64+0.64+0.64+0.64

= 2.56

Increase in homozygous frequency

Self fertilization of Gg = Gg* Gg

G(.2) g(0.8)
G (0.2) GG(0.04) Gg(0.16)
g(0.8) Gg(0.16) gg(0.64)

Frequencies in heterozygous self cross

GG = 0.04

Gg= 0.32

gg= 0.64

remains same

After many generation of self crossing homozygous individuals, homozygous frequency will increase and only homozygous will be present

In self crossing of heterozygous individuals, the frequency will remain same